OpenStudy (anonymous):

Which pair of ratios forms a proportion? 10/16 and 5/32 10/22 and 5/11 14/8 and 7/16 8/44 and 16/22

OpenStudy (mertsj):

A proportion is a statement that two ratios are equal.

OpenStudy (mertsj):

So for the first one write: \[\frac{10}{16}=\frac{5}{32}\]

OpenStudy (mertsj):

And then cross multiply to see if the statement is true

OpenStudy (mertsj):

Does 10x32=16x5?

OpenStudy (anonymous):

no it dosent

OpenStudy (mertsj):

So now, try the next one.

OpenStudy (anonymous):

nope that dont work either

OpenStudy (mertsj):

\[\frac{10}{22}=\frac{5}{11}\]

OpenStudy (mertsj):

What is 10 x 11?

OpenStudy (anonymous):

itss 110

OpenStudy (mertsj):

What is 5 x 22?

OpenStudy (anonymous):

110 (: OHHHHHH it is it!!]

OpenStudy (mertsj):

Yes.

OpenStudy (anonymous):

so the answer is 10/22=5/11

OpenStudy (mertsj):

The question says which pair forms a proportion and you should answer that 10/22 and 5/11 form a proportion.

OpenStudy (anonymous):

oh okay thank you(:

OpenStudy (mertsj):

yw